# Dual Alignment Maximin Optimization for Offline Model-based RL

### v2

@@ Appendix A Proofs in the Main Text
$=$ $(1-\gamma)\mathrm{E}_{s\sim\mu_{0},a\sim\pi}[Q(s,a)]$
Theorem A.7.
The optimal solution of maximin question (6)’s inner question $Q^{\star}(s,a)$ satisfies
$Q^{\star}(s,a)=r(s,a,s^{\prime})-\alpha\log\dfrac{\rho^{\pi}_{M}}{\rho^{\pi_{\beta}}_{T}}-\alpha f^{\prime}(\dfrac{\rho^{\pi}_{T}}{\rho^{\pi}_{M}})+\tau^{\pi}Q^{\star}(s,a)$ (7)
@@
Then the optimal solution of inner question is $y^{\star}(s,a,s^{\prime})=(f^{\prime}_{\star})^{-1}(\dfrac{\rho^{\pi}_{T}}{\rho^{\pi}_{M}})$. For any convex function, we have $(f^{\prime}_{\star})^{-1}(x)=f^{\prime}(x)$, so $y^{\star}(s,a,s^{\prime})=f^{\prime}(\dfrac{\rho^{\pi}_{T}}{\rho^{\pi}_{M}})$.
Recalling that we employ the variable substitution:$y(s,a,s^{\prime})=[r(s,a,s^{\prime})-\alpha\log\frac{\rho_{M}^{\pi}}{\rho_{T}^{\pi_{\beta}}}+\tau^{\pi}Q(s,a)-Q(s,a)]/\alpha$, so we have
$y^{\star}(s,a,s^{\prime})=\dfrac{r(s,a,s^{\prime})-\alpha\log\frac{\rho_{M}^{\pi}}{\rho_{T}^{\pi_{\beta}}}+\tau^{\pi}Q^{\star}(s,a)-Q^{\star}(s,a)}{\alpha}$
@@
$Q^{\star}(s,a)=r(s,a,s^{\prime})-\alpha\log\dfrac{\rho^{\pi}_{M}}{\rho^{\pi_{\beta}}_{T}}-\alpha f^{\prime}(\dfrac{\rho^{\pi}_{T}}{\rho^{\pi}_{M}})+\tau^{\pi}Q^{\star}(s,a)$
@@ Appendix B OOD Issues under Dynamics Model Integration
B.1 OOD Actions
The dynamics model generates data that can significantly deviate from the distribution of the offline dataset, leading to OOD actions. This triggers negative model exploitation (Levine et al., 2020), resulting in poor policy performance. Aligning the distributions of synthetic and offline data can effectively mitigate these issues. The model, $M$, is used to generate synthetic data for training, but it is essential to avoid taking OOD actions when rolling out policy in the model, to ensure accurate value estimation.
@@ B.2 OOD states
Appendix C Implementation Details
@@
C.1 Selection of f-divergence
As established in Theorem 5.3, the f-divergence must consistently exceed the KL-divergence to guarantee that the objective in Equation (2) serves as a lower bound for the standard RL objective $\mathcal{J}(\pi)$. Based on this, we specifically choose $f(x)=\frac{1}{3}(x-1)^{3}$, and its conjugate function is $f_{\star}(x)=\frac{2}{3}(x-1)^{\frac{3}{2}}.$ In principle, any choice of f-divergence is permissible, and we adhere to the selection made in the prior work (Luo et al., 2024b). Due to the tight schedule, we haven’t tried other choices of f-divergence. However, we suppose that the choice of f-divergence won’t enormously affect the effectiveness of DAMO, since f-divergence primarily serves as a measure of discrepancy.
Table 3: Hyperparameter of DAMO when optimizing policy
@@ C.1 Selection of f-divergence
C.2 Modification of actor training objective
When optimizing the actor network using objective (2), we observed that the substantial residual of the Bellman error often disrupts the optimization of the actor and leads to training instability. To stabilize the training process, we followed the implementation of previous DICE work (Mao et al., 2024a; Sikchi et al., 2023) and fixed the hyperparameter $\alpha$ to $1.0$ during the optimization of the actor network. The specific form of the actor training objective is as follows:
$\max\limits_{\pi}(1-\gamma)\mathrm{E}_{s\sim\mu_{0},a\sim\pi}[\tilde{Q}(s,a)]\\ +\mathrm{E}_{\rho_{M}^{\pi}}[f_{\star}(\Phi(s,a,s^{\prime}))]$
@@ C.2 Modification of actor training objective
C.3 Model, classifier and Policy Optimization
@@
Our model training approach is consistent with the methodology used in prior works (Yu et al., 2020; Sun et al., 2023). We train an ensemble of seven dynamics models and select the best five based on their validation prediction error from a held-out set containing 1000 transitions in the offline dataset $D_{R}$. Each model in the ensemble is a 4-layer feedforward neural network with 200 hidden units. We randomly choose one model from the best five models during model rollouts.
For classifier $h_{\psi}$, we implement the following approach: At each gradient step, we randomly sample data from both offline dataset $D_{R}$ and synthetic dataset $D_{M}$ with a fixed ratio $d$. This method ensures a balanced utilization of data from both DL and DG during training.
@@
Appendix D Experiment Setting
D.1 Benchamark
We evaluate our approach using the D4RL benchmark (Fu et al., 2020), focusing on the “v2” version of Gym tasks. For DICE-based methods (Algae-DICE, Opti-DICE, and O-DICE), we directly adopt the scores reported in their respective original papers. However, since the results for O-DICE on the random dataset are not provided in the original paper, we exclude them from our comparison. For offline model-based methods, we retrained MOPO (Yu et al., 2020) and MOBILE (Sun et al., 2023) on the “v2” Gym datasets. For the remaining methods, we directly report the scores provided in their original papers. All results are summarized in Table 2.
Table 4: Hyperparameters of DAMO used in the D4RL datasets.
@@ D.1 Benchamark
D.2 Hyperparameters
We provide a list of the hyperparameters that were tuned during our experiments. The detailed configurations are summarized in Table 4.
@@ D.2 Hyperparameters
The actor learning rate $l_{r}$ Since DAMO operates as a Maximin optimization problem and is susceptible to saddle points, we optimize the actor’s learning rate to achieve robust performance. The choice of learning rate plays a pivotal role in DAMO’s effectiveness. We tune the actor learning rate $l_{r}$ in the range of $\{1\times 10^{-5},5\times 10^{-5},1\times 10^{-4}\}$.
D.3 Computing Infrastructure

### v3

@@ A.2. Proof of Theorem 1
$=$ $(1-\gamma)\mathbb{E}_{s\sim\mu_{0},a\sim\pi}[Q(s,a)]$
In the first equality, we make a variable change of $y(s,a,s^{\prime})$. In the second equality, we explicitly formulate the operator $\tau^{\pi}$. In the third equality, we employ the proposition $\sum\limits_{s,a,s^{\prime}}\rho_{T}^{\pi}(s,a,s^{\prime})=1$. In the fourth equality, we express $\rho_{T}^{\pi}(s,a,s^{\prime})$ in the discounted sum of $\beta_{t}(s,a,s^{\prime})$. In the fifth equality, we employ the proposition $\beta_{t}(s,a,s^{\prime})=\beta_{t}(s,a,s^{\prime},a^{\prime})\pi(a^{\prime}|s^{\prime})$. In the sixth equality, we employ the proposition $\sum\limits_{s,a}\beta_{t}(s,a,s^{\prime},a^{\prime})=\beta(s^{\prime},a^{\prime})$.
For the second term in objective (5), we define $\Phi(s,a,s^{\prime})=\alpha y(s,a,s^{\prime})$, then objective (5) can be written as:
$\max\limits_{\pi}\min\limits_{Q(s,a)}(1-\gamma)\mathbb{E}_{s\sim\mu_{0},a\sim\pi}[Q(s,a)]+\alpha\mathbb{E}_{\rho_{M}^{\pi}}f_{\star}(\Phi(s,a,s^{\prime})/\alpha)$
We have now completed the proof of Theorem 1.
A.3. Proof of Theorem 2
Next, we present the proof of Theorem 2.
Theorem. 11.
If $Q^{\star}(s,a)$ is the optimal solution of inner minimization in maximin problem (6), then it satisfies
$Q^{\star}(s,a)=r(s,a,s^{\prime})-\alpha\log\dfrac{\rho^{\pi}_{M}}{\rho^{\pi_{\beta}}_{T}}-\alpha f^{\prime}(\dfrac{\rho^{\pi}_{T}}{\rho^{\pi}_{M}})+\mathcal{T}^{\pi}Q^{\star}(s,a)$ (7)
@@ A.3. Proof of Theorem 2
Therefore, the optimality condition is satisfied when the derivative of $y(s,a,s^{\prime})$ equals zero, and a closed-form solution of $y^{\star}(s,a,s^{\prime})$ can be obtained: $y^{\star}(s,a,s^{\prime})=(f^{\prime}_{\star})^{-1}(\dfrac{\rho^{\pi}_{T}}{\rho^{\pi}_{M}})$. Let $f(x)$ be a strictly convex function, then we have $(f^{\prime}_{\star})^{-1}(x)=f^{\prime}(x)$, thus $y^{\star}(s,a,s^{\prime})=f^{\prime}(\dfrac{\rho^{\pi}_{T}}{\rho^{\pi}_{M}})$.
Recalling that we employ the variable substitution: $y(s,a,s^{\prime})=[r(s,a,s^{\prime})-\alpha\log\frac{\rho_{M}^{\pi}}{\rho_{T}^{\pi_{\beta}}}+\mathcal{T}^{\pi}Q(s,a)-Q(s,a)]/\alpha$, so we have
$y^{\star}(s,a,s^{\prime})=\dfrac{r(s,a,s^{\prime})-\alpha\log\frac{\rho_{M}^{\pi}}{\rho_{T}^{\pi_{\beta}}}+\mathcal{T}^{\pi}Q^{\star}(s,a)-Q^{\star}(s,a)}{\alpha}$
@@
$Q^{\star}(s,a)=r(s,a,s^{\prime})-\alpha\log\dfrac{\rho^{\pi}_{M}}{\rho^{\pi_{\beta}}_{T}}-\alpha f^{\prime}(\dfrac{\rho^{\pi}_{T}}{\rho^{\pi}_{M}})+\mathcal{T}^{\pi}Q^{\star}(s,a)$
@@ Appendix B OOD Issues under Dynamics Model Integration
B.1. OOD Actions
The dynamics model generates data that can significantly deviate from the distribution of the offline dataset, leading to OOD actions. This triggers negative model exploitation Levine et al. (2020), resulting in poor policy performance. Aligning the distributions of synthetic and offline data can effectively mitigate these issues. The model, $M$, is used to generate synthetic data for training, but it is essential to avoid taking OOD actions when rolling out policy in the model, to ensure accurate value estimation.
@@ B.2. OOD states
Algorithm 1 DAMO
1:Offline dataset $\mathcal{D}_{R}$, dynamics model $M$, actor network $\pi_{\phi}$, critic network $Q_{\theta}$, classifier $h_{\psi}$.
2:Train the probabilistic dynamics model $M$ on $\mathcal{D}_{R}$
3:Initialize $\mathcal{D}_{M}\leftarrow\varnothing$
4:for each epoch do
5:  Generate synthetic rollouts by model $M$
6:  Add transitions to $\mathcal{D}_{M}$
7:  Update $h_{\psi}$ via Eq. (3)
8:  Calculate $\log(\rho_{M}^{\pi}/\rho_{T}^{\pi_{\beta}})$ via Eq. (4)
9:  Update $Q_{\theta}$ by minimizing Eq. (2)
10:  Update $\pi_{\phi}$ by maximizing Eq. (2)
11:end for
Appendix C Experiment Details
@@
C.1. Algorithm of DAMO
The complete algorithm of DAMO is summarized in Algorithm 1. Two replay buffers, $\mathcal{D}_{R}$ and $\mathcal{D}_{M}$, are used to store transitions from the offline dataset and model rollouts, respectively.
C.2. Selection of f-divergence
As established in Theorem 3, the f-divergence should consistently exceed the KL-divergence to guarantee that the objective in Equation (2) serves as a lower bound for the standard RL objective $\mathcal{J}(\pi)$. We adhere to the selection made in the prior work Luo et al. (2024a) and specifically choose $f(x)=\frac{1}{3}(x-1)^{3}$, its conjugate function is $f_{\star}(x)=\frac{2}{3}(x-1)^{\frac{3}{2}}.$ While for most values, $f(x)$ exceeds $x\log x$, in a small interval around $x=1$, $f(x)$ fails to statisfy the condition. In principle, it is feasible to choose a f-divergence satisfying the given condition, but due to the tight schedule, we haven’t tried other choices of f-divergence. However, we suppose that the choice of f-divergence won’t enormously affect the effectiveness of DAMO, since f-divergence primarily serves as a measure of discrepancy.
Table 5. Hyperparameter of DAMO when optimizing policy
@@ C.2. Selection of f-divergence
C.3. The inconsistent version of DAMO
Section 6.1.2 focuses on evaluating the effect of DAMO’s consistent policy improvement objective. Prior offline MBRL methods primarily concentrate on the design of conservative Q-functions. Once a conservative Q-function is obtained, these methods typically apply the SAC framework for policy improvement, i.e.,
$\max_{\pi}E_{s\sim D,a\sim\pi}Q^{\pi}(s,a)+\alpha\mathcal{H}(\pi).$
where $D$ denotes the offline dataset and $\mathcal{H}(\pi)$ denotes the entropy of the policy $\pi$.
In contrast, DAMO performs policy training directly based on the maximin formulation (2), where the policy improvement objective is consistent with the policy evaluation objective:
$\max_{\pi}\min_{Q}(1-\gamma)E_{s\sim\mu_{0},a\sim\pi}[Q(s,a)]+\alpha E_{\rho_{M}^{\pi}}[f_{\star}(\Phi(s,a,s^{\prime})/\alpha)].$
We refer to the variant of DAMO that applies the SAC-style policy improvement objective as the “inconsistent version.”
C.4. Modification of actor training objective
When optimizing the actor network using objective (2), we observed that the substantial residual of the Bellman error often disrupts the optimization of the actor and leads to training instability. To stabilize the training process, we followed the implementation of previous DICE work Mao et al. (2024b); Sikchi et al. (2023) and fixed the hyperparameter $\alpha$ to $1.0$ during the optimization of the actor network. The specific form of the actor training objective is as follows:
$\max\limits_{\pi}(1-\gamma)\mathbb{E}_{s\sim\mu_{0},a\sim\pi}[Q(s,a)]\\ +\mathbb{E}_{\rho_{M}^{\pi}}[f_{\star}(\Phi(s,a,s^{\prime}))]$
@@ C.4. Modification of actor training objective
C.5. The concrete setting of halcheetah-jump
Data collection. A SAC agent is trained in the halfcheetah task online for 1M environment steps. The reward function used in this phase is
$r(s,a)=\max\{v_{x},3\}-0.1\times\|a\|_{2}^{2},$
where $v_{x}$ denotes the velocity along the x-axis.
Reward shaping. After collecting the offline dataset, we relabel the reward function to:
$r(s,a)=\max\{v_{x},3\}-0.1\times\|a\|_{2}^{2}+15\times(z-initz),$
where $z$ denotes the z-position of the half-cheetah and init z denotes the initial z-position. As the observation of the original halfcheetah environment does not contain the information of init z, we add an extra dimension to the observation to inform the agent of init z.
Baselines. Three typical offline behavior-regularization methods are selected for comparison. BEAR Kumar et al. (2019)aims to constrain the policy’s actions to lie in the support of the behavioral distribution. BRAC Wu et al. (2019) is a family of algorithms that operate by penalizing the value function by some measure of discrepancy (KL divergence or MMD) between $\pi(\cdot|s)$ and $\pi^{B}(\cdot|s)$. TD3+BC Fujimoto and Gu (2021) introduces a behavior cloning regularization into the TD3 objective.
C.6. Model, classifier and Policy Optimization
@@
Our model training approach is consistent with the methodology used in prior works Yu et al. (2020); Sun et al. (2023). We train an ensemble of seven dynamics models and select the best five based on their validation prediction error from a held-out set containing 1000 transitions in the offline dataset $D_{R}$. Each model in the ensemble is a 4-layer feedforward neural network with 200 hidden units. We randomly choose one model from the best five models during model rollouts.
For classifier $h_{\psi}$, we implement the following approach: At each gradient step, we randomly sample data from both offline dataset $D_{R}$ and synthetic dataset $D_{M}$ with a fixed ratio $d$. This method ensures a balanced utilization of data from both $D_{R}$ and $D_{M}$ during training.
@@
Appendix D Additional Experiments
D.1. A quantitative analysis of Fig 1
When drawing Fig 1, we applied T-SNE to perform dimensionality reduction on the data and visualized the processed data to provide an intuitive understanding of data similarity. During the T-SNE process, we did not quantitatively measure the non-overlapping regions between datasets. To enable a more rigorous comparison, we computed the Wasserstein distance between the datasets. The results are in the Table 6. Experimental results indicate that all three methods achieve a good alignment between model data and offline data, but only DAMO successfully aligns model data with real data, which is consistent with our analysis in the paper. Comparing the directly computed Wasserstein distance with the Wasserstein distance after T-SNE dimensionality reduction, we observe that although the absolute values change, the relative ordering remains the same, confirming the validity of using T-SNE for visualization.
Table 6. Wasserstein distance between the offline, real, and model data. The first column clarifies the type of data, for example, Offline-Model means the Wasserstein distance between Offline data and Model data. ’w/ Tsne’ represents the usage of T-SNE to reduce the dimension of data and then calculate the Wasserstein distance.
Distance MOPO MOBILE DAMO (ours)
Offline-Model 0.24 0.13 0.16
Model-Real 0.39 0.43 0.27
Offline-Model(w/ T-SNE) 2.17 1.96 2.01
Model-Real(w/ T-SNE) 10.28 6.76 4.35
D.2. The performance in noisy environments
Table 7. Results on the noisy mujoco.
Task Name MOPO MOBILE DAMO (ours)
Halfcheetah-me 42.10 72.53 $74.51\pm 7.55$
Hopper-me 63.82 93.86 $95.36\pm 4.14$
Walker2d-me 108.69 114.90 $113.94\pm 1.15$
Noisy environment refers to a situation where random noise interferes with the agent’s state observation during testing, which is normal in real-world applications. Therefore, we have supplemented the following experiment in Mujoco to test whether DAMO is robust enough to demonstrate strong performance in noisy environment. We still use DAMO to train the policy on clean offline data, but during the testing phase, at each step transition, the environment introduces noise to the state observed by the agent with a certain probability. Under this setting, we tested the performance of DAMO, MOPO, and MOBILE, and the results are in the table 7. Both DAMO and MOBILE exhibit robust performance in noisy environments, effectively mitigating the impact of environmental noise on policy performance, whereas MOPO experiences a significant performance degradation. Notably, injecting noise in the Walker2d environment does not lead to a noticeable decline in policy performance, which may be attributed to the characteristics of the environment and task, warranting further investigation.
D.3. DAMO with different hyperparameters for dual alignment
Using different parameters for the two regularization terms in the DAMO can serve as a trick to improve algorithm performance. We conducted an experiment under the halfcheetah-medium-expert dataset to evaluate its impact. We set the coefficient of the term $D_{f}(\rho_{T}^{\pi}\|\rho_{M}^{\pi})$ in (1) to be $\beta$ and keep the coefficient of the term $\log(\rho_{M}^{\pi}/\rho_{T}^{\pi_{\beta}})$ to be $\alpha$. The training process is illustrated in Fig 6. This trick does not yield a significant effect on the Halfcheetah-Medium-Expert dataset. This may be attributed to two factors: DAMO itself is insensitive to the parameters $\alpha$ and $\beta$, and the Halfcheetah environment also exhibits low sensitivity to these parameters. The final scores from top to bottom are 103.1, 101.78, and 100.33.
Figure 6. fig-Setting different parameters for two penalization terms
Appendix E Experiment Setting
E.1. Benchamark
In the D4RL benchmark Fu et al. (2020), we focus on the “v2” version of Gym tasks. For DICE-based methods (Algae-DICE, Opti-DICE, and O-DICE), we directly adopt the scores reported in their respective original papers. However, since the results for O-DICE on the random dataset are not provided in the original paper, we exclude them from our comparison. For offline model-based methods, we retrained MOPO Yu et al. (2020) and MOBILE Sun et al. (2023) on the “v2” Gym datasets. For the remaining methods, we directly report the scores provided in their original papers. All results are summarized in Table 1.
In the NeoRL benchmark Qin et al. (2022), we report the performance of BC, CQL, and MOPO based on the original NeoRL paper, while the performance of TD3+BC, EDAC, and MOBILE is derived from the scores reported in the MOBILE Sun et al. (2023) paper.
Table 8. Hyperparameters of DAMO used in the D4RL datasets.
@@ E.1. Benchamark
E.2. Hyperparameters
We provide a list of the hyperparameters that were tuned during our experiments. The detailed configurations in the D4RL benchmark are summarized in Table 8. The detailed configurations in the NeoRL benchmark are summarized in Table 9. It should be noted that we considered two additional hyperparameters in the NeoRL benchmark, due to the narrow data coverage and noisy environments.
@@ E.2. Hyperparameters
The actor learning rate $l_{r}$. Since DAMO operates as a Maximin optimization problem and is susceptible to saddle points, we optimize the actor’s learning rate to achieve robust performance. The choice of learning rate plays a pivotal role in DAMO’s effectiveness. We tune the actor learning rate $l_{r}$ in the range of $\{1\times 10^{-5},5\times 10^{-5},1\times 10^{-4}\}$.
Offline data ratio. The limited coverage of the transition space leads to poor performance of the trained dynamics in some datasets. Consequently, we have to downweight the ratio of synthethic data used to optimize the objective for stabilizing the training process. We tune the offline data ratio within the range of $\{0.05,0.1,0.3\}$.
Auto-alpha. Similar to MOBILE, we forbid the automatic entropy tuning trick in some datasets.
Table 9. Hyperparameters of DAMO used in the NeoRL datasets.
Task Name $\alpha$ $K$ $l_{r}$ of actor Offline data ratio Auto-alpha
Halfcheetah-L 0.3 5 $1\times 10^{-4}$ 0.05 True
Hopper-L 1 5 $1\times 10^{-4}$ 0.05 True
Walker2d-L 2.5 1 $1\times 10^{-4}$ 0.1 True
Halfcheetah-M 0.5 3 $1\times 10^{-4}$ 0.05 True
Hopper-M 3 5 $5\times 10^{-5}$ 0.3 False
Walker2d-M 3.5 1 $1\times 10^{-5}$ 0.3 True
Halfcheetah-H 3 3 $1\times 10^{-4}$ 0.05 True
Hopper-H 5 5 $1\times 10^{-5}$ 0.05 False
Walker2d-H 3 3 $1\times 10^{-4}$ 0.3 True
E.3. Computing Infrastructure